OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10?

OpenStudy (ash2326):

Could you rearrange the equation such that, it's of the form \[y=mx+c\] ?

OpenStudy (anonymous):

Idk how to.

OpenStudy (ash2326):

divide whole equation by -2

OpenStudy (anonymous):

Even the 5x?

OpenStudy (anonymous):

rearrange it into the form of y = mx + c so 5x - 2y = -10 5x + 20 = 2y 2y = 5x + 20 y = 5/2x +20/2 y = 5/2x + 10 now the m = gradient or slope that you can read from the last line of the rearranged equation so what is it?

OpenStudy (ash2326):

yeah

OpenStudy (anonymous):

try it yourself and then check with mine.

OpenStudy (anonymous):

I got x=1?

OpenStudy (anonymous):

show me ur working out @Analiza?

OpenStudy (anonymous):

sorry to stike you down jay

OpenStudy (anonymous):

I got confused, how m I supose to divide -2y & 5x?

OpenStudy (anonymous):

can you understand the steps I did?

OpenStudy (anonymous):

5x - 2y = 10 so we want y to be on one side so we will have to +2y on both sides 5x -2y + 2y = 10 + 2y 5x = 10 + 2y now we have to minus 10 on both sides 5x - 10 = 10 -10 + 2y 5x - 10 = 2y now we are going to rewrite it 2y = 5x - 10 this is same as the previous step but just swapping them around y = 5/2 - 10/2 divide both sides by 2 y = 5/2x - 5

OpenStudy (anonymous):

lets try an easier sample, say 2x + y = 1, and rearrange into the form on y= mx + c

OpenStudy (anonymous):

slope is \[\frac{5}{2}\]

OpenStudy (anonymous):

yes, @laxkid we figured that out already

OpenStudy (anonymous):

I can see that you have trouble rearranging basic equations, so I think you should start by doing some exercises.